(3x+y-z)p+(x+y-z)q=2(z-y)
1
Expert's answer
2019-10-06T11:35:50-0400

The PDE has the form


"(3x + y - z)\frac{{\partial z}}{{\partial x}} + (x + y - z)\frac{{\partial z}}{{\partial y}} = 2(z - y)"

To solve this equation we need to find 2 independent solutions of characteristic system


"\frac{{dx}}{{3x + y - z}} = \frac{{dy}}{{x + y - z}} = \frac{{dz}}{{2(z - y)}}"

We can use the following property:


"\sum\limits_{k = 1}^n {\frac{{{a_k}}}{{{b_k}}}} = \frac{{\sum\limits_{k = 1}^n {{\lambda _k}{a_k}} }}{{\sum\limits_{k = 1}^n {{\lambda _k}{b_k}} }}"

First, choose "{\lambda _1} = - 1" , "{\lambda _2} = 3" and "{\lambda _3} = 1" and we get


"- dx + 3dy + dz = 0"

Then the first integral of the system is

"{\varphi _1} = - x + 3y + z"

Now let's choose "{\lambda _1} = 1" , "{\lambda _2} = - 1" , "{\lambda _3} = 1" and then "{\lambda _1} = 1" , "{\lambda _2} = 1" , "{\lambda _3} = -1" . Because in this case "\sum\limits_{k = 1}^n {{\lambda _k}{b_k}} \ne 0" these expressions will be equal so we get


"\frac{{dx - dy + dz}}{{2(x - y + z)}} = \frac{{dx + dy - dz}}{{4(x + y - z)}}"

"\frac{{d(x - y + z)}}{{x - y + z}} = \frac{1}{2}\frac{{d(x + y - z)}}{{x + y - z}}"

"\ln \left| {x - y + z} \right| = \ln \sqrt {\left| {x + y - z} \right|}"

Then the second integral is

"{\varphi _2} = x - y + z - \sqrt {x + y - z}"

Now we know 2 independent integrals of the characteristic system so we can write the solution of the PDE in the form


"F({\varphi _1};{\varphi _2}) = 0"

or explicitly


"F( - x + 3y + z;x - y + z - \sqrt {x + y - z} ) = 0"

where "F" is a smooth differentiable function.